# A Computational Social Simulation of Ageing and Care Accessibility in Italian Inner Areas

### v2

@@ Abstract
Ageing societies face increasing strain on formal and informal care systems, particularly in low-density mountainous municipalities where sparse services and steep terrain constrain access. This study presents a spatially explicit agent-based model that integrates a road-network GIS, synthetic populations derived through Iterative Proportional Fitting, and behavioural heterogeneity to examine how alternative service configurations shape accessibility and caregiver burden. The model, applied to Premeno (Piedmont, Italy), compares a baseline distribution of ambulatory services with a relocation scenario at Villa Bernocchi. System-level indicators (Caregiver Effort, Overwhelm, Hours Not Cared, Walkability) and micro-spatial metrics (Walkability, Detour Ratio, Proximity) are analysed across 40 batches and 50 stochastic replications per scenario. Results reveal aggregate neutrality but pronounced local redistribution of accessibility. Sensitivity analysis shows that spatial impedance dominates accessibility, whereas behavioural capacity modulates care effort. The findings illustrate hallmark properties of complex adaptive social systems-emergence, heterogeneity, and feedback-demonstrating how computational social simulation can illuminate policy trade-offs between spatial efficiency, social equity, and care sustainability in ageing territories.
keywords:
accessibility, agent-based modelling, ageing, caregiver burden, computational social science, GIS
@@ 1 Introduction
Demographic ageing is transforming the structure and functioning of contemporary societies. As life expectancy rises and family networks shrink, the capacity of public and informal care systems to guarantee accessible and sustainable assistance becomes increasingly uncertain. Peripheral and mountainous areas are particularly exposed: dispersed settlement patterns, steep topography, and scarce public transport combine to amplify inequalities in access to health and social services. Understanding these dynamics requires analytical frameworks able to capture both the spatial heterogeneity of environments and the interactive behaviour of the individuals embedded within them. Computational social science provides such a framework by combining data-rich representations of social environments with agent-based simulation of micro-level decisions and interactions [1, 2]. Within this perspective, agent-based models (ABMs) have been applied to domains ranging from urban mobility and spatial systems [3, 6] to health behaviour [9] and ageing policy [11]. Recent work has further situated ABM within population health frameworks [10]. Yet few studies integrate geographic information systems (GIS), demographic microdata, and survey-derived behavioural attributes in a single, reproducible simulation environment. This article presents a spatially explicit agent-based model (ABM) that couples a road network GIS with a behaviourally heterogeneous population of elder-caregiver dyads. The model integrates municipal indicators and survey-based parameters via synthetic population techniques to evaluate how ambulatory service reconfiguration affects both system-level performance and micro-spatial accessibility patterns. The empirical case focuses on the municipality of Premeno (Piedmont, Italy) and compares a baseline configuration (Scenario 1) with a relocation to Villa Bernocchi (Scenario 2). The study aims to quantify aggregate differences, reveal sub-group and neighbourhood heterogeneity, and identify policy-relevant trade-offs.
@@
Novel contribution. This study advances the field of computational social science by operationalising ageing-care accessibility as a complex adaptive social system. It unifies three methodological strands: (i) synthetic population generation through Iterative Proportional Fitting (IPF), (ii) GIS-embedded agent-based simulation of ageing and caregiving dynamics, and (iii) global sensitivity analysis for mechanism discovery. Together, these layers form a reproducible and data-rich framework that links the social-data synthesis tradition of population science with the mechanism-oriented exploration characteristic of computational social science. By integrating spatial, demographic, and behavioural dimensions, the model reveals how micro-level mobility and care-allocation behaviours give rise to macro-level accessibility and burden patterns, thereby supporting the evaluation of care sustainability in ageing territories.
@@ 2.1 Model overview
Input data Demographic and socioeconomic indicators from ISTAT; GIS data (OSM roads, Copernicus DEM); Structured caregiver interviews; Empirical mobility datasets; Literature-based behavioural parameters. See ESM 1 (§5)
@@
The model follows the updated ODD-D protocol [12], ensuring full transparency and reproducibility. A condensed checklist of its core components is reported in Table 1, while the complete protocol-including purpose, entities, state variables, process scheduling, and design concepts-is provided in the Electronic Supplementary Material (ESM 1). All stochastic processes (agent initialization, tie assignment, random choice) are controlled by a fixed random seed, ensuring reproducibility across runs. Initialization, warm-up, and randomisation procedures are detailed in ESM 1 (§5-6), including the Schruben-Singh-Tierney test for warm-up length and the hashing-based RNG seeding scheme. The model was developed in NetLogo 6.4 with Python extension for batch execution and sensitivity analysis.
@@ 2.2 Scenarios and experimental design
Figure 1: Study area: municipality of Premeno (Piedmont, Italy), an inner area characterised by low service density and steep morphology. The inset map (top right) shows its geographical position in northern Italy. Source: OpenStreetMap [osm].
@@ 2.3 Synthetic population and data integration
Two configurations are analysed for Premeno: Scenario 1 (baseline distribution of ambulatory services) and Scenario 2 (relocation to the Villa Bernocchi site). Each scenario was parameterised using empirical and literature-based constants summarised in Table 4 in the ESM (§5.1). Each scenario is executed in 40 experimental batches with 50 independent replications per batch (2,000 runs per scenario). Key Performance Indicators (KPIs) include: Caregiver Effort, Caregiver Overwhelm, Hours Not Cared, and a Walkability Index. Micro-level indices include Walkability (WKB), Effective Detour Ratio (EDR), and Household Proximity Index (HPI). Statistical comparisons employ Welch $t$-tests with multiplicity adjustment where applicable and cluster-robust checks as reported in the scenario analysis.
@@ 3 Sensitivity Analysis and Robustness
A two-stage global sensitivity analysis was conducted following Morris [7] and Saltelli and Annoni [8]. Stage 1 applied the Morris elementary-effects method (10 trajectories, $\Delta=0.5$) for qualitative screening; Stage 2 employed Sobol variance decomposition with Saltelli sampling ($n=512$ samples per parameter, N=2  [^0] N=2\,000 total model evaluations). Sensitivity analyses were implemented through an automated hybrid pipeline, wherein NetLogo-via its Python extension-triggered a command-line shell to launch a Jupyter notebook executing the SALib routines. This modular configuration enabled seamless coupling between model execution and parameter sampling while enforcing identical random seeds across replications to guarantee methodological reproducibility.
@@ 4.1 Emergent spatial patterns of accessibility
Under Scenario 1, walkability is highly heterogeneous within the municipal boundary. Values in the dedicated walkability analysis range approximately between 12 and 47 with a mean near 27, consistent with fragmented connectivity on steep terrain and distance to services. Under Scenario 2, the mean Walkability Index increases to about 0.34 and the variance decreases, indicating improved overall accessibility and greater spatial uniformity. However, micro-level indicators reveal a trade-off: route indirectness (EDR) increases and proximity (HPI) decreases for some central households. The relocation thus yields a more walkable environment with slightly less direct access for particular neighbourhoods (see Table 4).
4.2 Behavioural and social heterogeneity
Behavioural heterogeneity mediates how spatial change translates into caregiving burden. Higher-mobility dyads maintain acceptable accessibility even in peripheral areas, while mobility-limited dyads experience accelerated burden once care-capacity thresholds are reached. In the Premeno experiment, clusters of high-burden dyads coincide with lower walkability and higher dependency, evidencing differentiated outcomes across ageing stages. These patterns illustrate how individual-level variation and local morphology interact to produce uneven but interpretable social outcomes. Stage-wise results (Table 4) show that walkability (WKB) improves consistently across all stages, while the effective detour ratio (EDR) rises by roughly +0.9 to +1.1, signalling a spatial trade-off between usability and route efficiency. Proximity (HPI) exhibits small and often non-significant declines, reflecting marginal losses for central households.
4.3 System-level and micro-level effects
System-level KPI comparisons between Scenario 1 and Scenario 2 do not yield statistically significant differences after correction: caregiver effort changes are small, caregiver overwhelm declines marginally, walkability changes are negligible at the aggregate level, and hours not cared display a modest increase when using cluster-robust checks as reported in the scenario analysis (Table 3). By contrast, micro-spatial indicators reveal clear redistribution of accessibility: Scenario 2 enhances walkability and spatial uniformity while slightly increasing route detours (Table 4).
KPI Scenario 1 Scenario 2 $\boldsymbol{\Delta}$ (S2-S1) Significance
Caregiver Effort 0.16 0.17 +0.01 n.s. (Holm)
Hours Not Cared 1.86 2.16 +0.30 n.s. (Holm)
Caregiver Overwhelmed n.s. (see text)
Walkability (aggregate) $\approx 0$ n.s.; see micro-indices

### v3

@@ Abstract
Ageing societies face increasing strain on formal and informal care systems, particularly in low-density mountainous municipalities where sparse services and steep terrain constrain access. This study presents a spatially explicit agent-based model that integrates a road-network GIS, synthetic populations derived through Iterative Proportional Fitting, and behavioural heterogeneity to examine how alternative service configurations shape accessibility and caregiver burden. The model, applied to Premeno (Piedmont, Italy), compares a baseline distribution of ambulatory services with a relocation scenario at Villa Bernocchi. System-level indicators (Caregiver Effort, Overwhelmed Caregivers, Hours Not Cared, Walkability) and micro-spatial metrics (Walkability, Detour Ratio, Proximity) are analysed across 40 batches and 50 stochastic replications per scenario. Results reveal aggregate neutrality but pronounced local redistribution of accessibility. Sensitivity analysis shows that spatial impedance dominates accessibility, whereas behavioural capacity modulates care effort. The findings illustrate distinctive properties of complex adaptive social systems - emergence, heterogeneity, and feedback - demonstrating how computational social simulation can highlight policy trade-offs between spatial efficiency, social equity, and care sustainability in ageing territories.
Keywords: accessibility, agent-based modelling, ageing, caregiver burden, simulation modelling, GIS.
@@ 1 Introduction
Demographic ageing is transforming the structure and functioning of contemporary societies. As life expectancy rises and family networks shrink, the capacity of public and informal care systems to guarantee accessible and sustainable assistance becomes increasingly uncertain. Peripheral and mountainous areas are particularly exposed: dispersed settlement patterns, steep topography, and scarce public transport combine to amplify inequalities in access to health and social services. Understanding these dynamics requires analytical frameworks able to capture both the spatial heterogeneity of environments and the interactive behaviour of the individuals embedded within them. Computational social science provides such a framework by combining data-rich representations of social environments with agent-based simulation of micro-level decisions and interactions [4, 5]. Within this perspective, agent-based models (ABMs) have been applied to domains ranging from urban mobility and spatial systems [6, 13] to health behaviour [26] and ageing policy [2]. Recent work has further situated ABM within population health frameworks [24]. Yet few studies integrate geographic information systems (GIS), demographic microdata, and survey-derived behavioural attributes in a single, reproducible simulation environment. This article presents a spatially explicit agent-based model (ABM) that couples a road network GIS with a behaviourally heterogeneous population of elder-caregiver dyads. The model integrates municipal indicators and survey-based parameters via synthetic population techniques to evaluate how ambulatory service reconfiguration affects both system-level performance and micro-spatial accessibility patterns. The empirical case focuses on the municipality of Premeno (Piedmont, Italy) and compares a baseline configuration (Scenario 1) with a relocation of ambulatory services to Villa Bernocchi (Scenario 2), reflecting a policy option motivated by service rationalisation and cost-containment considerations. The study aims to quantify aggregate differences, reveal sub-group and neighbourhood heterogeneity, and identify policy-relevant trade-offs.
1.1 Positioning and Contribution
This study advances simulation-based analysis by operationalising ageing-care accessibility as a complex adaptive social system. Despite extensive research on ageing, accessibility, and service provision, a persistent gap remains in how care accessibility is conceptualised and analysed in sparsely populated and morphologically constrained territories. Much of the foundational and applied literature continues to operationalise accessibility through static indices, optimisation-based location models, or aggregate planning heuristics that prioritise average distance minimisation or coverage efficiency at the population level [12, 3, 6, 13]. While valuable for descriptive assessment and infrastructure benchmarking, these approaches typically abstract from individual-level behavioural constraints and implicitly assume representative users, fixed behavioural responses, and monotonic relationships between spatial efficiency and welfare outcomes [9, 20, 18, 22].
However, ageing-care systems exhibit the defining characteristics of complex socio-spatial systems, in which mobility limitations, caregiving capacity, and spatial structure interact non-linearly over time. Empirical and theoretical work in welfare geography, health services research, and computational social science shows that interventions improving aggregate accessibility indicators may simultaneously generate uneven redistributions of burden and opportunity across social groups and neighbourhoods [25, 1, 5, 4]. Such micro–macro decoupling is difficult to detect using aggregate or optimisation-based frameworks and remains underexplored in spatially explicit, agent-based analyses of care accessibility.
@@ 1.1 Positioning and Contribution
To address this gap, the present study unifies three methodological strands: (i) synthetic population generation through Iterative Proportional Fitting (IPF), (ii) GIS-embedded agent-based simulation of ageing and caregiving dynamics, and (iii) global sensitivity analysis for mechanism discovery. Together, these components form a reproducible and data-rich framework that links the social-data synthesis tradition of population science with the mechanism-oriented exploration characteristic of computational social science. By conceptualising accessibility and caregiver burden as emergent outcomes of decentralised, behaviourally heterogeneous interactions embedded in a real spatial network [14, 10], the model reveals a mechanism that is largely invisible to aggregate or optimisation-based approaches: cost-driven spatial reconfiguration can yield aggregate neutrality while redistributing accessibility and caregiver burden across neighbourhoods and social groups. The contribution of this work is therefore not a specific policy prescription, but a demonstration of how micro-level behavioural constraints and spatial structure decouple efficiency and equity at the system level, illustrating the analytical value of agent-based simulation for the study of ageing-care systems [26, 24, 11].
@@ 2 Methods
The empirical case is used as a controlled simulation environment designed to explore system behaviour under alternative service configurations, rather than as a basis for direct policy prescription or empirical forecasting. For clarity, we distinguish between different levels of the simulation design: a scenario refers to a structural configuration of the system (e.g., service location layout), a batch denotes a specific parameter setting within a scenario, and a run corresponds to a single stochastic replicate of a given batch.
@@ 2.1 Model overview
Input data Demographic and socioeconomic indicators from ISTAT; GIS data (OSM roads, Copernicus DEM); Structured caregiver interviews; Empirical mobility datasets; Literature-based behavioural parameters. See Supplementary Material (§5)
@@
The model follows the updated ODD-D protocol [11], ensuring full transparency and reproducibility. A condensed checklist of its core components is reported in Table 1, while the complete protocol-including purpose, entities, state variables, process scheduling, and design concepts-is provided in the Supplementary Material. All stochastic processes (agent initialization, tie assignment, random choice) are controlled by a fixed random seed, ensuring reproducibility across runs. Initialization, warm-up, and randomisation procedures are detailed in the Supplementary Material (§5-6), including the Schruben-Singh-Tierney test for warm-up length and the hashing-based RNG seeding scheme. The model was developed in NetLogo 6.4 with Python extension for batch execution and sensitivity analysis.
@@ 2.2 Scenarios and experimental design
Two configurations are analysed for Premeno: Scenario 1 (baseline distribution of ambulatory services) and Scenario 2 (relocation to the Villa Bernocchi site). Each scenario was parameterised using empirical and literature-based constants summarised in Table S4 in the Supplementary Material (§5.1). Each scenario is executed in 40 experimental batches with 50 independent replications per batch (2,000 runs per scenario). Key Performance Indicators (KPIs) include: Caregiver Effort Index (CEI), Caregiver Overwhelmed (CO), Hours Not Cared (HNC), and a Walkability Index (WKB); see Table 3. Micro-level indices include Walkability (WKB), Effective Detour Ratio (EDR), and Household Proximity Index (HPI). Statistical comparisons employ Welch $t$-tests with multiplicity adjustment where applicable and cluster-robust checks as reported in the scenario analysis. In this study, accessibility denotes potential spatial opportunity to reach services, operationalised via the Walkability Index ($\mathrm{WKB}$), whereas realised care outcomes are captured by caregiver- and dyad-level indicators, including the Caregiver Effort Index ($\mathrm{CEI}$), Hours Not Cared ($\mathrm{HNC}$), and the Caregiver Overwhelmed indicator ($\mathrm{CO}$).
Figure 1: Study area: municipality of Premeno (Piedmont, Italy), an inner area characterised by low service density and steep morphology. The inset map (top right) shows its geographical position in northern Italy. Source: OpenStreetMap [21].
@@ 3 Sensitivity Analysis and Robustness
A two-stage global sensitivity analysis was conducted following Morris [19] and Saltelli & Annoni [23]. Stage 1 applied the Morris elementary-effects method (10 trajectories, $\Delta=0.5$) for qualitative screening; Stage 2 employed Sobol variance decomposition with Saltelli sampling ($n=512$ samples per parameter, N=2  [^0] N=2\,000 total model evaluations). Sensitivity analyses were implemented through an automated hybrid pipeline, wherein NetLogo-via its Python extension-triggered a command-line shell to launch a Jupyter notebook executing the SALib routines. This modular configuration enabled seamless coupling between model execution and parameter sampling while enforcing identical random seeds across replications to guarantee methodological reproducibility.
@@
Table 3: Key performance indicators (KPIs) and notation used throughout the study. Units refer to the model’s internal time and distance conventions (see Supplementary Material for implementation details).
Symbol Name Unit Interpretation Aggregation level
$\mathrm{WKB}$ Walkability Index score (0–50) Potential spatial accessibility to services via the road network under mobility constraints node / household / dyad; summarised to municipality
$\mathrm{CEI}$ Caregiver Effort Index effort score (0–1) Care effort required from the informal caregiver to meet elder needs given accessibility and constraints dyad; summarised by stage / municipality
$\mathrm{CO}$ Caregiver Overwhelmed number Count of the caregivers unable to provide more care; proxy for overload risk dyad; summarised by stage / municipality
$\mathrm{HNC}$ Hours Not Cared hours per period Unmet care demand (hours of required care not delivered) under current constraints dyad; summarised by stage / municipality
  • Notes. All KPIs are internally normalized before aggregation. $\mathrm{WKB}$, $\mathrm{CEI}$, and $\mathrm{CO}$ are dimensionless indices derived from time- and network-based quantities via monotone transformations, while $\mathrm{HNC}$ is expressed in physical time units but may be rescaled for comparability. Aggregated values are computed as population- or dyad-weighted means unless stated otherwise. Formal definitions and implementation details for $\mathrm{WKB}$, $\mathrm{CEI}$, $\mathrm{CO}$, and $\mathrm{HNC}$ are provided in the Supplementary Material (SM, Sections S2–S4).
Table 4: Micro-level spatial and network indices used for mechanism analysis. Indices are computed locally using network-based distances and routing.
Symbol Name Unit Interpretation Aggregation level
$\mathrm{WKB}$ Walkability Index score (0–50) Local potential to reach essential services on foot via the road network under mobility constraints node / household / dyad
$\mathrm{EDR}$ Effective Detour Ratio ratio ($\geq 1$) Inefficiency of network-based walking routes relative to straight-line distance node / household / dyad
$\mathrm{HPI}$ Household Proximity Index score (dimensionless) Network-based proximity of a household to nearby households or support nodes household / dyad
  • Notes. All indices use network-based distances; Euclidean distance is used only as a reference for $\mathrm{EDR}$. $\mathrm{WKB}$ values are internally normalised and discretised for reporting. Formal definitions and computational details are provided in the Supplementary Material (SM, Sections S2–S3).
3.1 Alternative Methodological Approaches
Several alternative methodological approaches could, in principle, be used to study ageing-care accessibility in spatially constrained territories. These approaches are briefly discussed here to clarify the rationale for adopting an agent-based modelling framework.
First, static accessibility metrics and GIS-based indicators—such as cumulative opportunity measures, gravity-based indices, or isochrone analyses—are widely used to evaluate spatial access to services [12, 9, 22]. While effective for benchmarking infrastructure coverage and identifying underserved areas, these methods treat accessibility as a fixed property of space and population averages. They do not capture how individual mobility limitations, caregiving capacity, or scheduling constraints interact dynamically over time, nor how accessibility gains translate into realised care under heterogeneous behavioural conditions [18, 20].
Second, optimisation-based location and allocation models—such as maximal coverage, $p$-median, or network flow formulations—can be employed to design service configurations that minimise average distance or maximise population coverage under resource constraints [3]. Although valuable for planning and normative assessment, such models typically assume representative users and deterministic demand, and they optimise system-level objectives by construction. Consequently, they are ill-suited to analysing behavioural adaptation, distributional effects, or the emergence of unintended inequalities following spatial interventions, particularly in ageing-care systems characterised by heterogeneous constraints [6, 13].
Third, regression-based and quasi-experimental approaches using observational data have been applied to assess associations between accessibility measures and health or care outcomes [1, 25]. However, these methods require strong assumptions about exogeneity and stable unit treatment values, and they are limited in their ability to explore counterfactual spatial configurations. This limitation is especially pronounced in small or peripheral municipalities, where data sparsity and the absence of natural experiments constrain causal inference.
In contrast, the agent-based approach adopted in this study is specifically designed to address these limitations. By representing accessibility and caregiver burden as emergent outcomes of decentralised interactions between heterogeneous agents embedded in a real spatial network, the model enables the exploration of non-linear dynamics, micro–macro decoupling, and distributional effects that are not accessible through aggregate, optimisation-based, or purely statistical frameworks [14, 5, 4]. The aim is therefore not to replace existing methods, but to complement them by providing a mechanism-oriented perspective on ageing-care accessibility in complex territorial settings, consistent with established principles of agent-based modelling and computational social science [10, 11, 26, 24].
@@ 4 Results and Discussion
This study contributes to computational social science by demonstrating how ageing-care accessibility emerges from the interaction between spatial morphology, behavioural heterogeneity, and decentralised decision-making. Rather than producing uniform improvements, spatial reconfiguration of services generates compensatory dynamics: aggregate indicators remain largely unchanged, while accessibility and caregiver burden are redistributed across neighbourhoods and social groups.
From a theoretical perspective, this result illustrates a key property of complex adaptive social systems: efficiency-oriented interventions can be system-neutral yet socially redistributive. The model shows that spatial efficiency and social equity are not monotonically aligned, as behavioural constraints condition how spatial opportunities translate into realised accessibility. This mechanism is largely invisible to representative- agent models or static accessibility metrics, but becomes evident when micro-level heterogeneity and spatial networks are modelled explicitly.
The study also highlights the importance of multi-scale evaluation. While system-level indicators suggest stability, micro-spatial and stage-stratified analyses reveal differentiated effects, underscoring the risk of relying exclusively on aggregate metrics in policy assessment. Agent-based simulation thus serves not as a forecasting tool, but as an explanatory device for identifying structural trade-offs and unintended consequences of policy interventions.
@@ 4.1 Emergent spatial patterns of accessibility
A central finding of the simulation is the coexistence of aggregate neutrality with pronounced local and group-level redistribution of accessibility and caregiving burden. While system-level indicators show limited change across scenarios, micro-spatial and stage-stratified metrics reveal differentiated gains and losses, indicating a divergence between macro-level stability and micro-level reallocation. Under Scenario 1, walkability is highly heterogeneous within the municipal boundary. Values in the dedicated walkability analysis range approximately between 12 and 47 with a mean near 27, consistent with fragmented connectivity on steep terrain and distance to services. Under Scenario 2, the mean Walkability Index increases to about 34 and the variance decreases, indicating improved overall accessibility and greater spatial uniformity. However, micro-level indicators reveal a trade-off: route indirectness (EDR) increases and proximity (HPI) decreases for some central households. The relocation thus yields a more walkable environment with slightly less direct access for particular neighbourhoods (see Table 5). Such micro–macro divergence is a characteristic feature of agent-based representations of social systems, where aggregate stability can coexist with substantial local redistribution driven by heterogeneous constraints and interactions [5, 4].
@@
4.2 Behavioural and social heterogeneity
The redistribution observed across scenarios is not solely a spatial effect but is strongly conditioned by behavioural heterogeneity among elder–caregiver dyads. Differences in mobility, caregiving capacity, income, and support networks shape how spatial accessibility translates into realised care outcomes. Behavioural heterogeneity mediates how spatial change translates into caregiving burden. Higher-mobility dyads maintain acceptable accessibility even in peripheral areas, while mobility-limited dyads experience accelerated burden once care-capacity thresholds are reached. In the Premeno experiment, clusters of high-burden dyads coincide with lower walkability and higher dependency, evidencing differentiated outcomes across ageing stages. These patterns illustrate how individual-level variation and local morphology interact to produce uneven but interpretable social outcomes. Stage-wise results (Table 5) show that walkability (WKB) improves consistently across all stages, while the effective detour ratio (EDR) rises by roughly +0.9 to +1.1, signalling a spatial trade-off between usability and route efficiency. Proximity (HPI) exhibits small and often non-significant declines, reflecting marginal losses for central households. These findings reinforce the role of behavioural heterogeneity as a conditioning mechanism that filters spatial opportunity, a dynamic that cannot be represented by representative-agent or purely spatial accessibility models [10, 26].
4.3 System-level and micro-level effects
The relocation scenario considered in this study reflects a cost-containment and service-rationalisation logic frequently discussed in small-municipality health planning. Rather than assuming accessibility improvement as an objective, the scenario provides a test case for examining the distributive consequences of efficiency-driven spatial reconfiguration. System-level KPI comparisons between Scenario 1 and Scenario 2 do not yield statistically significant differences after correction: caregiver effort changes are small, caregiver overwhelm declines marginally, walkability changes are negligible at the aggregate level, and hours not cared display a modest increase when using cluster-robust checks as reported in the scenario analysis (Table 6). By contrast, micro-spatial indicators reveal clear redistribution of accessibility: Scenario 2 enhances walkability and spatial uniformity while slightly increasing route detours (Table 5). From this perspective, the simulation highlights how efficiency-oriented interventions may reallocate disadvantage without producing aggregate welfare gains, a pattern consistent with broader discussions of resilience and unintended effects in public health and welfare systems [7, 27].
KPI Scenario 1 Scenario 2 $\boldsymbol{\Delta}$ (S2-S1) Significance
Caregiver Effort 0.16 0.17 +0.01 n.s. (Holm)
Hours Not Cared 1.86 2.16 +0.30 n.s. (Holm)
Caregiver Overwhelmed n.s. (see text)
Walkability (aggregate) $\approx 0$ n.s.; see micro-indices
As with all agent-based models, the present simulation is intended as an explanatory and exploratory device rather than a predictive instrument. Its scope is bounded by stylised behavioural rules, exogenously specified institutional configurations, and simplified representations of health and care trajectories. This divergence between macro-neutrality and micro-variation illustrates that spatial interventions reallocate relative advantage rather than uniformly improving welfare. Such heterogeneous effects, typical of complex adaptive systems, highlight the need for multi-scale evaluation when interpreting policy outcomes. Explicitly acknowledging these boundary conditions clarifies the class of questions the model is designed to address and aligns with established best practices for transparent and reproducible agent-based modelling [11, 5]. Accordingly, validation is approached in an exploratory and structural sense, focusing on the internal coherence of mechanisms, the plausibility of agent behaviours, and the robustness of observed patterns under parameter variation. Rather than seeking point-wise empirical prediction, the model is intended to assess whether hypothesised interactions between spatial structure, behavioural constraints, and service configuration are sufficient to generate the observed qualitative patterns of accessibility and caregiver burden. Global sensitivity analysis is therefore used as a primary validation instrument to identify influential parameters and to distinguish structural effects from artefacts of parameterisation. This approach is consistent with established practices in agent-based and simulation modelling when the aim is to explore mechanisms and trade-offs in complex social systems.